OpenStudy (goformit100):

An error 2% in excess is made while measuring the side of a square. The percentage of error in the calculated area of the square is: A. 2% B. 2.02% C. 4% D. 4.04%

OpenStudy (goformit100):

@mayankdevnani

OpenStudy (goformit100):

HELP

OpenStudy (mayankdevnani):

i think we need 100 cm side of square!!!!

OpenStudy (mayankdevnani):

so 2% of 100=102 cm!!! got it @goformit100

OpenStudy (kropot72):

D. is correct.

OpenStudy (goformit100):

yes i got it

OpenStudy (mayankdevnani):

it means that area1=100*100=100^2. correct!!!!

OpenStudy (goformit100):

you both are simply confusing me

OpenStudy (mayankdevnani):

got it!!! @goformit100

OpenStudy (goformit100):

yes got that :) THAKS

OpenStudy (mayankdevnani):

welcome!!!!

OpenStudy (kropot72):

\[1.02^{2}=1.0404\] Error is therefore \[\frac{1.0404-1}{1}\times 100=4.04\ percent\]